# Agentic-R1: Distilled Dual-Strategy Reasoning

### v1

@@ 2 Related Work
While prior efforts have integrated external tools into language models Gao et al. (2023); Schick et al. (2023); Nakano et al. (2022), they are often specialized to either non-math domains or are confined to shorter reasoning chains. Concurrently, the paradigm of long chain-of-thought (long-CoT) reasoning or inference-time compute has demonstrated significant improvements Guo et al. (2025); Jaech et al. (2024). However, these approaches can be difficult to control and may suffer from ‘overthinking’, particularly when applied to tool-use scenarios Cuadron et al. (2025). Recent works have combined tool use with long CoT reasoning Feng et al. (2025); Song et al. (2025), but these are often applied to different domains or rely on reinforcement learning, which can be less stable than our proposed distillation method. To the best of our knowledge, DualDistill is the first framework to employ distillation with trajectory composition from two heterogeneous teacher models, one specializing in agentic tool-use and the other in pure textual reasoning, creating a unified student model capable of adaptively leveraging both strategies. See Appendix B for a more detailed discussion.
@@ 3.4 Self-Distillation
To address it, we introduce self-distillation to help the student further refine its strategy selection based on its capabilities and the given problem. Our self-distillation process involves the student model generating candidate solutions, with teacher models providing verification or corrections as auxiliary supervision. The process reinforces effective strategies and corrects suboptimal ones. Specifically, given a training set and the student policy finetuned on distillation data from teachers , we sample trajectories for each problem :
@@ 4 Experiments
| Model | Budget | DeepMath-L | Combinatorics300 | MATH500 | AIME | AMC | avg. |
| Qwen2.5-7B-Instruct (w/o tool) | S L | 17.2 17.5 | 21.8 21.8 | 75.1 75.2 | 8.0 8.0 | 42.9 42.9 | 33.0 33.1 |
| Qwen2.5-7B-Instruct (w/ tool) | S L | 34.7 34.7 | 28.9 28.9 | 70.2 70.2 | 14.7 14.7 | 51.1 51.1 | 39.9 39.9 |
| DeepSeek-R1-Distill-7B | S L | 34.7 56.3 | 34.7 44.5 | 83.1 89.2 | 23.3 40.7 | 61.2 84.8 | 47.4 63.1 |
| Agentic-R1-7B (ours) | S L | 37.0 59.3 | 36.9 49.4 | 80.0 82.4 | 28.0 40.7 | 64.3 82.2 | 49.3 62.8 |
| Agentic-R1-7B-SD (ours) | S L | 40.0 65.3 | 38.2 52.0 | 82.5 93.3 | 27.3 40.7 | 66.3 85.8 | 50.9 67.4 |
@@ 5 Conclusion
While our approach demonstrates strong overall performance, several limitations remain for future work. First, the transition words connecting different strategies within composed trajectories are currently designed manually. As a result, the output produced by the trained student model can occasionally lack naturalness and fluidity, especially when switching between strategies. Developing methods for more coherent and automatic transitions of strategies could further enhance the readability and overall quality of the content generated by the student model.
@@ A.2 Training Configuration
For finetuning the student model Agentic-R1 on teacher distilled trajectories, we use 4 A6000 GPUs for a total of 12.7 hours. The model is trained for 4 epochs using the AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of . We set the maximum context length to 16,384 tokens for teacher distillation and 8,192 for self-distillation, and discard all training examples that exceed this limit.
@@ A.3.1 Problem Filtering Heuristics
Difficulty Under Constraints: We use a baseline text reasoning-only model, Deepseek-R1-Distill-7B, with a limited context length (4096 tokens). Problems unsolvable under the constraint with one trail are deemed more difficult and suitable for agentic strategies.
@@ A.3.2 Dataset Scale
After running the two teachers on the filtered subset and compositing the trajectories, the final distilled dataset contains examples. The detailed count for each correctness category is listed in Tab. 3.
@@ A.4.2 Trajectory Composition Implementation
To transform multiturn agentic trajectories from OpenHands logs into a suitable training format, we extract content from the log fields labeled ‘thought’, ‘code’, and ‘final thought’ along with their associated executor feedback if any. Each extracted content is then enclosed within distinct resource tags—<think></think>, <code></code>, <answer></answer> or <executor></executor>—and concatenated sequentially. For reasoning trajectories from Deepseek-R1, we specifically apply the <answer></answer> tag to the content outside the long CoT part (i.e., beyond the <think></think> segment).
@@ A.5 Qualitative Example
We observed that Agentic-R1 shows several promising behaviors: (1) The model initially adopts tool-aided reasoning, but results in wrong results after several attempts, and then the model automatically switches to text reasoning and finally derives the correct answer (Fig. 6); (2) The model initially tries to apply text reasoning for a combinatorial problem, and then changes to tool-aided reasoning to reduce computational complexity (Fig. 7).
@@ A.6 Inference Details
For all evaluation experiments, we use the VLLM framework Kwon et al. (2023) to enable fast inference via prefix caching, which significantly accelerates multi-turn tool calling. In the tool-augmented setting, the language model is allowed to invoke a Python executor up to times per problem, with each execution capped at seconds. During inference, whenever the model outputs the special token </code>, the generation process is temporarily paused, the preceding code block is executed, and the resulting feedback is appended to the ongoing generation enclosed with <executor> </executor> before resuming inference. Although tool execution introduces up to seconds of additional runtime per query, this cost is relatively small compared to the time-intensive pure text reasoning process, which can take several minutes to reach a conclusion using Deepseek-R1-Distill-7B on 2A6000 GPUs. Additionally, the prompt template is listed in Fig. 2.
@@
Exact match: The grader extracts the final non-think block: a code result (‘<executor>…</executor>’) or a direct answer (‘<ans>…</ans>’), and compares it with the gold answer.
@@ Appendix B Related Work
Integrating external tools into the language model chain-of-thoughts has substantially improved the accuracy of numerical and factual tasks. Early program-aided methods such as PaL Gao et al. (2023) and PoT Chen et al. (2023) demonstrated significant gains by converting reasoning steps into executable programs, delegating precise computations to code interpreters. Other lines of work, including WebGPT Nakano et al. (2022) and ReAct Yao et al. (2023), introduced agent-like reasoning frameworks that interleave tool invocation (e.g., web searches or API calls) within multi-step reasoning. Toolformer Schick et al. (2023) further generalized this approach by training language models to self-supervise API calls on various tasks such as arithmetic, translation, and retrieval. However, unlike DualDistill, these methods typically rely on short chain-of-thoughts primarily using prompting or heuristic-based tool invocation, lacking mechanisms for automatically balancing pure long reasoning against tool use based on task complexity.
@@
Recent approaches have highlighted significant performance improvements by scaling inference-time chain-of-thought length. OpenAI’s O1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025) used outcome-driven reinforcement learning to generate extensive reasoning trajectories, substantially outperforming shorter-CoT baselines on complex math and reasoning benchmarks. Similarly, S1 Muennighoff et al. (2025) and L1 Aggarwal and Welleck (2025) demonstrated scaling curves showcasing a log-linear relationship between performance and inference compute. Empirical evidence supports that an increase in inference computation can often yield more cost-effective gains than an increase in model size alone Wu et al. (2025). Nonetheless, long-CoT models frequently encounter overthinking, i.e, generating overly long reasoning that leads to redundant or incorrect outcomes, especially in tool use scenarios, in a phenomenon known as the reasoning-action dilemma Cuadron et al. (2025). Our work addresses these issues by teaching a student model when to switch between internal reasoning and tool-based execution adaptively.
@@
Recently, some works have explored the idea of combining long-form reasoning with explicit tool invocation. R1-Searcher Song et al. (2025) and Search-R1 Jin et al. (2025) introduced reinforcement-learning-based retrieval policies within reasoning loops, achieving substantial performance improvements on open-domain question-answering tasks. However, unlike these methods, DualDistill is specifically tailored for math tasks. Similarly, ReTool Feng et al. (2025) trained a reasoning model with tool use for math tasks. However, unlike these approaches that rely on expensive and unstable reinforcement learning techniques, DualDistill is a simple distillation approach, leading to a more data-efficient and practical training setup.
@@
A growing line of work explores methods for language models to automatically improve from their own generated outputs or feedback, a paradigm often referred to as self-taught learning or self-distillation. Early approaches in this area explored self-taught learning with pseudo-labels, where a model generates rationales or answers on unlabeled data and subsequently fine-tunes itself on these predictions Zelikman et al. (2022); Wang et al. (2022). Shen et al. (2022) introduces self-distillation by using the outputs of the previous mini-batch as soft targets, improving output consistency without external teachers. In the LLM era, fine-tuning LLMs on self-generated trajectories with a rule-based verifier has been applied in many domains, including coding tasks Jiang et al. (2024), formal proving Lin et al. (2024), math tasks Kumar et al. (2024), and others. A concurrent work Zheng et al. (2025) enables self-evolving reasoning across multiple modalities and fuses them during inference.

### v2

@@ 2 Related Work
Although prior efforts have integrated external tools into language models Gao et al. (2023); Schick et al. (2023); Nakano et al. (2022), they are often specialized to non-math domains or are confined to shorter reasoning chains. Concurrently, the paradigm of long chain-of-thought (long-CoT) reasoning or inference-time compute has demonstrated significant improvements Guo et al. (2025); Jaech et al. (2024). However, these approaches can be difficult to control and may suffer from ‘overthinking’, particularly when applied to tool-use scenarios Cuadron et al. (2025). Recent works have combined tool use with long-CoT reasoning Feng et al. (2025); Song et al. (2025), but these are often applied to different domains or rely on reinforcement learning, which can be less stable than our proposed distillation method. To the best of our knowledge, DualDistill is the first framework to employ distillation with trajectory composition from two heterogeneous teacher models, one specializing in agentic tool-use and the other in pure textual reasoning, creating a unified student model capable of adaptively leveraging both strategies. See Appendix B for a more detailed discussion.
@@ 3.4 Self-Distillation
To address it, we introduce self-distillation to help the student further refine its strategy selection based on its capabilities and the given problem. Our self-distillation process involves the student model generating candidate solutions, with teacher models providing verification or corrections as auxiliary supervision. The process reinforces effective strategies and corrects suboptimal ones. Specifically, given a training set and the student policy fine-tuned on distillation data from teachers , we sample trajectories for each problem :
@@ 4 Experiments
| Model | Budget | DeepMath-L | Combinatorics300 | MATH500 | AIME | AMC | avg. | ||||||||||||||
| Qwen2.5-7B-Instruct (w/o tool) |
| Qwen2.5-7B-Instruct (w/ tool) |
| DeepSeek-R1-Distill-7B |
| Agentic-R1-7B (ours) |
| Agentic-R1-7B-SD (ours) |
@@ 5 Conclusion
While our approach demonstrates strong overall performance, several limitations remain for future work. First, the transition words connecting different strategies within composed trajectories are currently designed manually. As a result, the output produced by the trained student model can occasionally lack naturalness and fluidity, especially when switching between strategies. Moreover, the student model after self-distillation may exhibit repetitive behavior. Developing methods for more coherent and automatic transitions between strategies could further enhance the readability and overall quality of the content generated by the student model.
@@
Acknowledgments
This work was supported in part by the National Science Foundation under Grant Nos. DMS-2434614 and DMS-2502281.
@@ A.2 Training Configuration
For fine-tuning the student model Agentic-R1 on teacher distilled trajectories, we use 4 A6000 GPUs for a total of 12.7 hours. The model is trained for 4 epochs using the AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of . We set the maximum context length to 16,384 tokens for teacher distillation and 8,192 for self-distillation, and discard all training examples that exceed this limit.
@@ A.3.1 Problem Filtering Heuristics
Difficulty Under Constraints: We use a baseline text reasoning-only model, DeepSeek-R1-Distill-7B, with a limited context length (4096 tokens). Problems unsolvable under the constraint with one trial are deemed more difficult and suitable for agentic strategies.
@@ A.3.2 Dataset Scale
After running the two teachers on the filtered subset and composing the trajectories, the final distilled dataset contains examples. The detailed count for each correctness category is listed in Tab. 3.
@@ A.4.2 Trajectory Composition Implementation
To transform multi-turn agentic trajectories from OpenHands logs into a suitable training format, we extract content from the log fields labeled ‘thought’, ‘code’, and ‘final thought’ along with their associated executor feedback, if any. Each extracted content is then enclosed within distinct resource tags—<think></think>, <code></code>, <answer></answer> or <executor></executor>—and concatenated sequentially. For reasoning trajectories from DeepSeek-R1, we specifically apply the <answer></answer> tag to the content outside the long CoT part (i.e., beyond the <think></think> segment).
@@ A.5 Qualitative Example
We observed that Agentic-R1 shows several promising behaviors: (1) The model initially adopts tool-aided reasoning, but yields incorrect outputs after several attempts, and then the model automatically switches to text reasoning and finally derives the correct answer (Fig. 6); (2) The model initially tries to apply text reasoning for a combinatorial problem, and then changes to tool-aided reasoning to reduce computational complexity (Fig. 7).
@@ A.6 Inference Details
For all evaluation experiments, we use the VLLM framework Kwon et al. (2023) to enable fast inference via prefix caching, which significantly accelerates multi-turn tool calls. In the tool-augmented setting, the language model is allowed to invoke a Python executor up to times per problem, with each execution capped at seconds. During inference, whenever the model outputs the special token </code>, the generation process is temporarily paused, the preceding code block is executed, and the resulting feedback is appended to the ongoing generation enclosed with <executor> </executor> before resuming inference. Although tool execution introduces up to seconds of additional runtime per query in the worst case, this cost is relatively small compared to the time-intensive pure text reasoning process, which can take several minutes to reach a conclusion using DeepSeek-R1-Distill-7B on 2A6000 GPUs. Additionally, the prompt template is listed in Fig. 2.
@@
Exact match: The grader extracts the final non-think block: a code result (‘<executor>…</executor>’) or a direct answer (‘<answer>…</answer>’), and compares it with the gold answer.
@@ Appendix B Related Work
Integrating external tools into the language model chain-of-thought (CoT) has substantially improved the accuracy of numerical and factual tasks. Early program-aided methods, such as PaL Gao et al. (2023) and PoT Chen et al. (2023), demonstrated significant gains by converting reasoning steps into executable programs, thereby delegating precise computations to code interpreters. Other lines of work, including WebGPT Nakano et al. (2022) and ReAct Yao et al. (2023), introduced agent-like reasoning frameworks that interleave tool invocation (e.g., web searches or API calls) within multi-step reasoning. Toolformer Schick et al. (2023) further generalized this approach by training language models to self-supervise API calls on various tasks such as arithmetic, translation, and retrieval. START Li et al. (2025b) and CoRT Li et al. (2025a) use hint-based prompting to activate tool use behavior, followed by rejection sampling fine-tuning (RFT) for self-improvement. However, unlike DualDistill, these methods typically focus on short-CoT, or primarily use prompting or heuristic-based tool invocation, lacking mechanisms to automatically balance long-term reasoning against tool use based on task complexity.
@@
Recent approaches have highlighted significant performance improvements by scaling the length of the inference time chain-of-thought (CoT). GPT-o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025) used outcome-driven reinforcement learning to generate extensive reasoning trajectories, substantially outperforming short-CoT baselines in complex math and reasoning benchmarks. Similarly, S1 Muennighoff et al. (2025) and L1 Aggarwal and Welleck (2025) demonstrated scaling curves showcasing a log-linear relationship between performance and inference compute. Empirical evidence supports that an increase in inference computation can often yield more cost-effective gains than an increase in model size alone Wu et al. (2025). However, long-CoT models frequently encounter overthinking, which is the generation of overly long reasoning that leads to redundant or incorrect outcomes, especially in tool use scenarios, a phenomenon known as the reasoning-action dilemma Cuadron et al. (2025). Our work addresses these issues by teaching a student model when to switch between internal reasoning and tool-based execution adaptively.
@@
Recently, some works have explored the idea of combining long-form reasoning with explicit tool invocation. R1-Searcher Song et al. (2025) and Search-R1 Jin et al. (2025) introduced reinforcement-learning-based retrieval policies within reasoning loops, achieving substantial performance improvements in open-domain question-answering tasks. However, unlike these methods, DualDistill is specifically tailored for math tasks. Similarly, ReTool Feng et al. (2025) trained a reasoning model with tool use for math tasks. However, unlike these approaches that rely on expensive and unstable reinforcement learning techniques, DualDistill is a simple distillation approach, leading to a more data-efficient and practical training setup.
@@
A growing line of work explores methods for language models to automatically improve from their own generated outputs or feedback, a paradigm often referred to as self-taught learning or self-distillation. Early approaches in this area explored self-taught learning with pseudo-labels, where a model generates rationales or answers on unlabeled data and subsequently fine-tunes itself on these predictions Zelikman et al. (2022); Wang et al. (2022). Shen et al. (2022) introduces self-distillation using the output of the previous mini-batch as soft targets, improving the consistency of the output without external teachers. In the LLM era, fine-tuning LLMs on self-generated trajectories with a rule-based verifier has been applied in many domains, including coding tasks Jiang et al. (2024), theorem proving Lin et al. (2024), math tasks Kumar et al. (2024), and others. A concurrent work Zheng et al. (2025) enables self-evolving reasoning across multiple modalities and fuses them during inference.